Nickblade10:

PLZ HELP ME: Show all work to identify the asymptotes and state the end behavior of the function f of x is equal to 6x divided by the quantity of x minus 36 end quantity.

Nickblade10:

umm:

Given: \[f(x)\frac{ 6x }{ x-36 }\] Apply rule of (example):\[(a)=a\] So it'd be this for our function: \[(x)=x\] \[=fx \frac{ 6x }{ x-36 }\] Apply fraction rule(example): \[a \times \frac{ b }{ c }=\frac{ a \times b }{ c }\] Our function will look like: \[=\frac{ fx \times 6x }{ x-36 }\] Apply exponent rule (example):\[aa=a^2\] Our function: \[xx=x^2\] \[=\frac{ fx^2 \times 6 }{ x-36 }\] Leading us with our final answer: \[=\frac{ 6fx^2 }{ x-36 }\]